# The Spatial Confounding Environment

### v2

@@ 1 Introduction
To address this challenge, we present SpaCE: The Spatial Confounding Environment, a comprehensive set of software tools and datasets for evaluating machine-learning methods to address spatial confounding in causal inference. SpaCE datasets comprise real treatment and confounder data from publicly available sources commonly used in environmental health, social science, economics, and climate science studies, among other domains. Core to our approach is generating realistic semi-synthetic outcomes approximating true outcomes of interest using state-of-art machine-learning ensembles (Erickson et al., 2020) paired with cross-validation procedures for spatial data (Roberts et al., 2017) and spatial correlation modeling. In each dataset, there is a set of masked spatially-varying covariates inducing spatial confounding, and categorized by degree of spatial smoothness. SpaCE datasets contain spatial information in an adjacency graph and/or geographical coordinates that methods can utilize. The diversity of SpaCE datasets, illustrated by Fig. 1, spans multiple treatment types (binary, continuous), spatial formats (graph, coordinates), and evaluation metrics (average causal effects, dose-response curves, or counterfactual prediction at the unit level). In addition to our SpaCE Python package, we provide tools to reproduce existing datasets and generate new ones from a user’s data. [^1] The SpaCE source code is available at https://github.com/NSAPH-Projects/space
@@ 3 SpaCE: The Spatial Confounding Environment
SpaceEnv
Type     #SpaceDatasets
Social Vulnerability and Welfare
census tract Texas 16 6,828 21,585
cdcsvi_limteng_hburdic_cont
continuous 12
cdcsvi_nohsdp_poverty_cont
continuous 11
cdcsvi_nohsdp_poverty_disc
binary 11
Heat Exposure and Wildfires
census tract California 22 8,616 26,695
climate_relhum_wfsmoke_cont
continuous 8
climate_wfsmoke_minrty_disc
binary 10
Air Pollution and Mortality
county USA 34 3,109 9,237
healthd_dmgrcs_mortality_disc
binary 9
healthd_hhinco_mortality_cont
continuous 10
healthd_pollutn_mortality_cont
continuous 9
Welfare and Elections
county USA 45 3,109 9,237
county_educatn_election_cont
continuous 14
county_phyactiv_lifexpcy_cont
continuous 16
county_dmgrcs_election_disc
binary 14
$\text{PM}_{2.5}$ Components
$1\times 1$ km USA 6 3,882,956 7,647,552
pm25_hires_no3_pm_cont
continuous 4
pm25_hires_so2_pm_cont
continuous 4
Socioeconomic Status and Broadband Usage
zip code USA 16 30,383 174,345
zcta_income_broadband_disc
discrete 5
zcta_age_broadband_cont
continuous 4
@@ 4 Details about the Data Generation Pipeline
[…]
@@
Training data
${\bm{X}}^{\text{obs}},{\bm{A}}$ and $\tilde{{\bm{Y}}}$
Counterfactuals
$\tilde{{\bm{Y}}}^{a}$ for each treatment value $a\in{\mathbb{A}}$, computed with Eq. 1. For continuous treatments, a discretization of $|{\mathbb{A}}|=100$ values is used
Graph and coordinates
A list of edges and coordinate matrix
Treatment type
An indicator if $|{\mathbb{A}}|=2$ (binary) or $|{\mathbb{A}}|=100$ (continuous)
Spatial smoothness score
Auto-correlation of ${\bm{X}}^{\text{miss}}$
Confounding score
A metric of the degree of confounding induced by ${\bm{X}}^{\text{miss}}$. .
@@
Table 3: Supported causal estimation tasks and metrics. $\tau_{*}$ denotes a causal effect and $\hat{\tau}_{*}$ is an estimate, where *** is the name of the task. For the last column, $\hat{Y}_{s}^{a}$ is an estimate of the counterfactual $\tilde{Y}_{s}^{a}$ for each location $s$ and treatment $a$. For continuous treatment types, we assume $|{\mathbb{A}}|=100$. The metrics are standardized by the standard deviation of $\tilde{Y}_{s}$, denoted $\sigma_{y}$.
Causal Effect
Treatment type
Metric
average treatment effect (ate)
binary
absolute bias (bias)
$\tau_{\text{ate}}=|{\mathbb{S}}|^{-1}\sum_{s\in{\mathbb{S}}}(\tilde{Y}_{s}^{1}-\tilde{Y}_{s}^{0})$
$\sigma_{y}^{-1}|\hat{\tau}_{\text{ate}}-\tau_{\text{ate}}|$
exposure-response function (erf)
continuous
root mean integrated-squared error ($\sqrt{\textsc{mise}}$)
$\forall a\in{\mathbb{A}}:\tau_{\text{erf}}(a)=|{\mathbb{S}}|^{-1}\sum_{s\in{\mathbb{S}}}\tilde{Y}_{s}^{a}$
$\sigma_{y}^{-1}\sqrt{|{\mathbb{A}}|^{-1}\sum_{a\in{\mathbb{A}}}(\hat{\tau}_{\text{erf}}(a)-\tau_{\text{erf}}(a))^{2}}$
individualized treatment effects (ite) $\forall a\in{\mathbb{A}},s\in{\mathbb{S}}:\tilde{Y}_{s}^{a}$
binary and continuous
precision at estimating heterogeneous effects (pehe) $\sigma_{y}^{-1}\sqrt{|{\mathbb{A}}|^{-1}{\mathbb{S}}|^{-1}\sum_{s\in{\mathbb{S}},a\in{\mathbb{A}}}(\tilde{Y}_{s}^{a}-\hat{Y}_{s}^{a})^{2}}$
@@ Appendix A Data collections
[…]
@@ Appendix B SpaceEnv Generation Details
The codebase for generating an environment is provided separately from the main package at https://github.com/NSAPH-Projects/space-data. Users can use this codebase to generate new SpaceEnvs from existing or new data collections using the steps outlined below. See the repository’s documentation for details.
@@
Each DataCollection is documented in the package repository and documentation. We include a description of the data sources, purpose, exploratory data analysis, and other relevant information. The contents of the documentation is inspired by data nutrition labels (Stoyanovich & Howe, 2019). This documentation can be accessed (anonymized) at https://github.com/NSAPH-Projects/space-data.
@@ Appendix D Hyper-parameter Tuning and Additional details of benchmarks
Model
Iterations
Tuning Metric
Value
gcnn
2,500
Dimension of the hidden layers (hidden_dim)
16 or 64
Number of hidden layers (hidden_layers)
1 or 2
Weight decay (weight_decay)
1e-6 to 1e-1
spatial+
2,500
lam_t
loguniform between 1e-5 and 1.0
lam_y
loguniform between 1e-5 and 1.0
spatial
2,500
lam
loguniform between 1e-5 and 1.0
dapsm
N/A
propensity_score_penalty_value
Choice among [0.001, 0.01, 0.1, 1.0]
propensity_score_penalty_type
Choice between l1 and l2
spatial_weight
Uniform between 0.0 and 0.1
caliper_type
daps
matching_algorithm
optimal
@@ Appendix E Licenses of New and Reused Assets
Resource name
License
Python package SpaCE  [^2] https://github.com/NSAPH-Projects/space
MIT
Model-training software repository space-data  [^3] https://github.com/NSAPH-Projects/space-data
MIT
Benchmark data collection hosted at Harvard Dataverse
CC0 1.0
@@
Data source
Reference
License
Synthetic Medicare Data for Environmental Health Studies
https://doi.org/10.7910/DVN/L7YF2G
CC0 1.0
NOAA
https://www.nauticalcharts.noaa.gov/data/data-licensing.html
CC0 1.0
CDC
https://wonder.cdc.gov/DataUse.html
Public domain
U.S. Census Bureau
https://ask.census.gov/prweb/PRServletCustom?pyActivity=pyMobileSnapStart&ArticleID=KCP-4928
Public domain
The Bureau of Labor Statistics (BLS)
https://www.bls.gov/opub/copyright-information.html
Public domain
PRISM
https://prism.oregonstate.edu/terms/
Open for non-commercial purposes
Police shooting by the Washington Post
https://github.com/washingtonpost/data-police-shootings
CC BY-NC-SA 4.0
US Broadband Data
https://github.com/microsoft/USBroadbandUsagePercentages
Open Use of Data Agreement v1.0
@@
https://sedac.ciesin.columbia.edu/data/set/aqdh-pm2-5-concentrations-contiguous-us-1-km-2000-2016
Public domain
@@
https://www.ciesin.columbia.edu/data/aqdh/pm25component-EC-NH4-NO3-OC-SO4-2000-2019/
Public domain
@@ Appendix G Q&A
GgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v3

@@ 1 Introduction
To address this challenge, we present SpaCE: The Spatial Confounding Environment, a comprehensive set of software tools and datasets for evaluating machine-learning methods to address spatial confounding in causal inference. SpaCE datasets comprise real treatment and confounder data from publicly available sources commonly used in environmental health, social science, economics, and climate science studies, among other domains. Core to our approach is generating realistic semi-synthetic outcomes approximating true outcomes of interest using state-of-art machine-learning ensembles (Erickson et al., 2020) paired with cross-validation procedures for spatial data (Roberts et al., 2017) and spatial correlation modeling. In each dataset, there is a set of masked spatially-varying covariates inducing spatial confounding, and categorized by degree of spatial smoothness. SpaCE datasets contain spatial information in an adjacency graph and/or geographical coordinates that methods can utilize. The diversity of SpaCE datasets, illustrated by Fig. 1, spans multiple treatment types (binary, continuous), spatial formats (graph, coordinates), and evaluation metrics (average causal effects, dose-response curves, or counterfactual prediction at the unit level). In addition to our SpaCE Python package, we provide tools to reproduce existing datasets and generate new ones from a user’s data. [^1] The SpaCE source code is available at https://anonymous.4open.science/r/space-BC93
@@ 3 SpaCE: The Spatial Confounding Environment
SpaceEnv Type #SpaceDatasets
Social Vulnerability and Welfare census tract Texas 16 6,828 21,585
cdcsvi_limteng_hburdic_cont continuous 12
cdcsvi_nohsdp_poverty_cont continuous 11
cdcsvi_nohsdp_poverty_disc binary 11
Heat Exposure and Wildfires census tract California 22 8,616 26,695
climate_relhum_wfsmoke_cont continuous 8
climate_wfsmoke_minrty_disc binary 10
Air Pollution and Mortality county USA 34 3,109 9,237
healthd_dmgrcs_mortality_disc binary 9
healthd_hhinco_mortality_cont continuous 10
healthd_pollutn_mortality_cont continuous 9
Welfare and Elections county USA 45 3,109 9,237
county_educatn_election_cont continuous 14
county_phyactiv_lifexpcy_cont continuous 16
county_dmgrcs_election_disc binary 14
$\text{PM}_{2.5}$ Components $1\times 1$ km USA 6 3,882,956 7,647,552
pm25_hires_no3_pm_cont continuous 4
pm25_hires_so2_pm_cont continuous 4
Socioeconomic Status and Broadband Usage zip code USA 16 30,383 174,345
zcta_income_broadband_disc discrete 5
zcta_age_broadband_cont continuous 4
@@ 4 Details about the Data Generation Pipeline
  • Residual sampling. For (b) and (c), residual sampling, we use a Gaussian Markov Random Field (GMRF) from a spatial graph, a convenient and scalable spatial process (Rue & Held, 2005; Tec et al., 2019). More precisely, we sample the synthetic residuals as ${\bm{R}}\sim_{\text{iid}}\text{MultivariateNormal}({\bm{0}},\hat{\lambda}({\mathbb{D}}-\hat{\rho}{\mathbb{A}}{\mathbb{D}})^{-1})$, where ${\mathbb{A}}$ is the adjacency matrix of the spatial graph; ${\mathbb{D}}$ is a diagonal matrix with the number of neighbors of each location; $\hat{\rho}$ determines the correlation between an observation and its neighbors, learned from the true residuals (the errors of the predictive model in Step 1); and $\hat{\lambda}$ is chosen to match the variance of the true residuals exactly. A key technical element in our implementation was sampling from a large Gaussian random field using matrix factorizations for sparse matrices (Chen et al., 2008), enabling scalability to massive spatial graphs (such as in the $\text{PM}_{2.5}$ Components DataCollection).
@@
Training data ${\bm{X}}^{\text{obs}},{\bm{A}}$ and $\tilde{{\bm{Y}}}$
Counterfactuals $\tilde{{\bm{Y}}}^{a}$ for each treatment value $a\in{\mathbb{A}}$, computed with Eq. 1. For continuous treatments, a discretization of $|{\mathbb{A}}|=100$ values is used
Graph and coordinates A list of edges and coordinate matrix
Treatment type An indicator if $|{\mathbb{A}}|=2$ (binary) or $|{\mathbb{A}}|=100$ (continuous)
Spatial smoothness score Auto-correlation of ${\bm{X}}^{\text{miss}}$
Confounding score A metric of the degree of confounding induced by ${\bm{X}}^{\text{miss}}$. .
@@
Table 3: Supported causal estimation tasks and metrics. $\tau_{*}$ denotes a causal effect and $\hat{\tau}_{*}$ is an estimate, where * is the name of the task. For the last column, $\hat{Y}_{s}^{a}$ is an estimate of the counterfactual $\tilde{Y}_{s}^{a}$ for each location $s$ and treatment $a$. For continuous treatment types, we assume $|{\mathbb{A}}|=100$. The metrics are standardized by the standard deviation of $\tilde{Y}_{s}$, denoted $\sigma_{y}$.
Causal Effect Treatment type Metric
average treatment effect (ate) binary absolute bias (bias)
$\tau_{\text{ate}}=|{\mathbb{S}}|^{-1}\sum_{s\in{\mathbb{S}}}(\tilde{Y}_{s}^{1}-\tilde{Y}_{s}^{0})$ $\sigma_{y}^{-1}|\hat{\tau}_{\text{ate}}-\tau_{\text{ate}}|$
exposure-response function (erf) continuous root mean integrated-squared error ($\sqrt{\textsc{mise}}$)
$\forall a\in{\mathbb{A}}:\tau_{\text{erf}}(a)=|{\mathbb{S}}|^{-1}\sum_{s\in{\mathbb{S}}}\tilde{Y}_{s}^{a}$ $\sigma_{y}^{-1}\sqrt{|{\mathbb{A}}|^{-1}\sum_{a\in{\mathbb{A}}}(\hat{\tau}_{\text{erf}}(a)-\tau_{\text{erf}}(a))^{2}}$
individualized treatment effects (ite) $\forall a\in{\mathbb{A}},s\in{\mathbb{S}}:\tilde{Y}_{s}^{a}$ binary and continuous precision at estimating heterogeneous effects (pehe) $\sigma_{y}^{-1}\sqrt{|{\mathbb{A}}|^{-1}{\mathbb{S}}|^{-1}\sum_{s\in{\mathbb{S}},a\in{\mathbb{A}}}(\tilde{Y}_{s}^{a}-\hat{Y}_{s}^{a})^{2}}$
@@ Appendix A Data collections
  • The $\text{PM}_{2.5}$ Components collection combines the high resolution total $\text{PM}_{2.5}$ dataset of Di et al. (2019) with the $\text{PM}_{2.5}$ composition data from Amini et al. (2022), both available at high spatial resolution of $1\times 1$ km grid. We use the datasets corresponding to the annual average for 2000.
@@ Appendix B SpaceEnv Generation Details
The codebase for generating an environment is provided separately from the main package at https://anonymous.4open.science/r/space-data-0C73. Users can use this codebase to generate new SpaceEnvs from existing or new data collections using the steps outlined below. See the repository’s documentation for details.
@@
Each DataCollection is documented in the package repository and documentation. We include a description of the data sources, purpose, exploratory data analysis, and other relevant information. The contents of the documentation is inspired by data nutrition labels (Stoyanovich & Howe, 2019). This documentation can be accessed (anonymized) at https://anonymous.4open.science/r/space-data-0C73.
@@ Appendix D Hyper-parameter Tuning and Additional details of benchmarks
Model Iterations Tuning Metric Value
gcnn 2,500 Dimension of the hidden layers (hidden_dim) 16 or 64
Number of hidden layers (hidden_layers) 1 or 2
Weight decay (weight_decay) 1e-6 to 1e-1
spatial+ 2,500 lam_t loguniform between 1e-5 and 1.0
lam_y loguniform between 1e-5 and 1.0
spatial 2,500 lam loguniform between 1e-5 and 1.0
dapsm N/A propensity_score_penalty_value Choice among [0.001, 0.01, 0.1, 1.0]
propensity_score_penalty_type Choice between l1 and l2
spatial_weight Uniform between 0.0 and 0.1
caliper_type daps
matching_algorithm optimal
@@ Appendix E Licenses of New and Reused Assets
Resource name License
Python package SpaCE  [^2] https://anonymous.4open.science/r/space-BC93 MIT
Model-training software repository space-data  [^3] https://anonymous.4open.science/r/space-data-0C73 MIT
Benchmark data collection hosted at Harvard Dataverse CC0 1.0
@@
Data source Reference License
Synthetic Medicare Data for Environmental Health Studies https://doi.org/10.7910/DVN/L7YF2G CC0 1.0
NOAA https://www.nauticalcharts.noaa.gov/data/data-licensing.html CC0 1.0
CDC https://wonder.cdc.gov/DataUse.html Public domain
U.S. Census Bureau https://ask.census.gov/prweb/PRServletCustom?pyActivity=pyMobileSnapStart&ArticleID=KCP-4928 Public domain
The Bureau of Labor Statistics (BLS) https://www.bls.gov/opub/copyright-information.html Public domain
PRISM https://prism.oregonstate.edu/terms/ Open for non-commercial purposes
Police shooting by the Washington Post https://github.com/washingtonpost/data-police-shootings CC BY-NC-SA 4.0
US Broadband Data https://github.com/microsoft/USBroadbandUsagePercentages Open Use of Data Agreement v1.0